# Skeleton Recall Loss for Connectivity Conserving and Resource Efficient Segmentation of Thin Tubular Structures

### v1

@@ Abstract
Accurately segmenting thin tubular structures, such as vessels, nerves, roads or concrete cracks, is a crucial task in computer vision. Standard deep learning-based segmentation loss functions, such as Dice or Cross-Entropy, focus on volumetric overlap, often at the expense of preserving structural connectivity or topology. This can lead to segmentation errors that adversely affect downstream tasks, including flow calculation, navigation, and structural inspection. Although current topology-focused losses mark an improvement, they introduce significant computational and memory overheads. This is particularly relevant for 3D data, rendering these losses infeasible for larger volumes as well as increasingly important multi-class segmentation problems. To mitigate this, we propose a novel Skeleton Recall Loss, which effectively addresses these challenges by circumventing intensive GPU-based calculations with inexpensive CPU operations. It demonstrates overall superior performance to current state-of-the-art approaches on five public datasets for topology-preserving segmentation, while substantially reducing computational overheads by more than . In doing so, we introduce the first multi-class capable loss function for thin structure segmentation, excelling in both efficiency and efficacy for topology-preservation. Our code is available to the community, providing a foundation for further advancements at: [will be shared later]
@@ 3.2.2 Soft Recall on Tubed Skeleton
Following the extraction of our tubed skeleton, we incentivize the network to include as much of this skeleton as possible as part of the prediction. This is performed simply by using a soft recall loss , in addition to any existing generic loss used by the network (for example, Dice Loss, Cross Entropy Loss, etc.), as seen in Algorithm 2. This vastly improves the connectivity of thin curvilinear structures predicted by a network trained using this loss (Sec. 5.1). Additionally, Skeleton Recall Loss is computationally inexpensive in comparison to the use of a differentiable skeletonization, requiring only fractionally more GPU memory and additional time during training (Sec. 5.3). This facilitates training multi-class segmentation problems as opposed to current differentiable skeleton methods which incur infeasible overheads (Sec. 5.3.2).
@@ 6 Conclusion
Acknowledgements
The present contribution is supported by the Helmholtz Association under the joint research school "HIDSS4Health – Helmholtz Information and Data Science School for Health."

### v2

@@ Abstract
Accurately segmenting thin tubular structures, such as vessels, nerves, roads or concrete cracks, is a crucial task in computer vision. Standard deep learning-based segmentation loss functions, such as Dice or Cross-Entropy, focus on volumetric overlap, often at the expense of preserving structural connectivity or topology. This can lead to segmentation errors that adversely affect downstream tasks, including flow calculation, navigation, and structural inspection. Although current topology-focused losses mark an improvement, they introduce significant computational and memory overheads. This is particularly relevant for 3D data, rendering these losses infeasible for larger volumes as well as increasingly important multi-class segmentation problems. To mitigate this, we propose a novel Skeleton Recall Loss, which effectively addresses these challenges by circumventing intensive GPU-based calculations with inexpensive CPU operations. It demonstrates overall superior performance to current state-of-the-art approaches on five public datasets for topology-preserving segmentation, while substantially reducing computational overheads by more than . In doing so, we introduce the first multi-class capable loss function for thin structure segmentation, excelling in both efficiency and efficacy for topology-preservation. Our code is available to the community, providing a foundation for further advancements, at: https://github.com/MIC-DKFZ/Skeleton-Recall.
@@ 3.2.2 Soft Recall on Tubed Skeleton
Following the extraction of our tubed skeleton, we incentivize the network to include as much of this skeleton as possible as part of the prediction. This is performed simply by using a soft recall loss (Eq. 1), in addition to any existing generic loss used by the network (for example, Dice Loss, Cross Entropy Loss, etc.).
This vastly improves the connectivity of thin curvilinear structures predicted by a network trained using this loss (Sec. 5.1). Additionally, Skeleton Recall Loss is computationally inexpensive in comparison to the use of a differentiable skeletonization, requiring only fractionally more GPU memory and additional time during training (Sec. 5.3). This facilitates training multi-class segmentation problems as opposed to current differentiable skeleton methods which incur infeasible overheads (Sec. 5.3.2).
@@ 6 Conclusion
Acknowledgement
The present contribution is supported by the Helmholtz Association under the joint research school "HIDSS4Health – Helmholtz Information and Data Science School for Health". This work was partly funded by Helmholtz Imaging (HI), a platform of the Helmholtz Incubator on Information and Data Science. PV is funded through an Else Kröner Clinician Scientist Endowed Professorship by the Else Kröner Fresenius Foundation (reference number: 2022_EKCS.17).